# From Static to Adaptive Defense: Federated Multi-Agent Deep Reinforcement Learning-Driven Moving Target Defense Against DoS Attacks in UAV Swarm Networks

### v1

@@ III-C Moving Target Defense Mechanisms
Figure 2: An overview of the proposed MTD-based mitigation mechanisms, including (a) leader switching, (b) route mutation, and (c) frequency hopping.
@@
To address these challenges, we propose a suite of lightweight and coordinated MTD mechanisms specifically tailored for UAV swarm networks. These mechanisms are designed to leverage the inherent dynamicity and flexibility of UAVs, enabling effective defense with minimal resource consumption and latency. Fig. 2 provides an overview of the proposed MTD-based mitigation mechanisms.
@@ III-C1 Leader Switching
Leader switching designates a new UAV as the swarm leader when the current leader is persistently targeted or isolated by attacks. It is important to note that leader switching is a logical (virtual) role reassignment that can be executed rapidly and with minimal overhead and does not require any physical movement within the swarm. Let $\mathbb{L}^{\mathcal{N}}(t)$ denote the swarm leader at time $t$. The system can switch to a new leader chosen from the set of eligible UAVs after the leader switching occurs:
@@
Here, $i$ cannot be 00 (the GCS) or the current leader and $\tau^{\mathbb{L}}_{\text{exec}}$ denotes the execution time that the leader switching mechanism takes. After the switch, the new leader takes over the responsibility of receiving control commands from the GCS and disseminating them to the follower UAVs through the flying Ad-Hoc network (FANET).
@@ III-C2 Route Mutation
Usually, the command is transferred from the GCS to the leader, and is then forwarded to the follower UAVs. However, when a direct communication link is blocked by an attack, we exploit route mutation that selects an alternative relay UAV to forward messages. For example, when the communication link between the current leader $\mathbb{L}^{\mathcal{N}}(t)=i$ and a follower UAV $j$ is under attack, then the relay UAV $r\in\mathcal{V}\setminus\{0,i,j\}$ should satisfy the following rules:
@@
This operation dynamically reconfigures the communication graph in real time, allowing the command to be forwarded through the relay UAV and successfully delivered to the follower, thus mitigating the impact of the targeted link attack, i.e., $E^{\mathcal{L}}_{ij}(t+\tau^{\mathbb{R}}_{\text{exec}})=0$, where $\tau^{\mathbb{R}}_{\text{exec}}$ is the time for updating the route table. When there is no ongoing attack or when the attack fails to produce a substantial impact, the system can further optimize network performance by restoring the relay-based communication path back to a direct link. This reduces communication latency and improves overall efficiency without compromising the defense capability.
@@ III-D Problem Formulation
For the UAV swarm network, its objective is to defend against the malicious DoS attacker. The UAVs need to choose the proper defensive actions to maximize the security performance, maintain the formation, and ensure communication connectivity while minimizing the defense consumption in a distributed setting. Therefore, we formulate the optimization problem for the UAV swarm network security as follows:
@@
where $|\mathcal{E}(t)|$ is the cardinality of this set and $C^{i}(t)$ is the defense cost of UAV $i$ at each time slot. Constraints C1 restrict the bound of the distance between UAVs, the attack effectiveness, and the communication links. Constraint C2 ensures that the heading angle of each UAV is within the range $[0,2\pi)$. The patrol speed of each UAV to be within the range of $[r\omega,v_{\max}]$ is enforced by Constraint C3, where $r$ is the formation radius and $\omega$ is the angular speed. Constraint C4 ensures that all UAVs operate on the same frequency channel at any time step, which is essential for maintaining communication connectivity. Finally, the attack duration and reconnaissance period are non-negative integers, allowing for a well-defined attack cycle, as delineated in constraint C5.
@@ III-E Optimization Analysis
It is worth noting that the multi-objective optimization problem above might not be solved by using traditional methods since obtaining accurate knowledge of the system dynamics and the attacker’s behavior is impractical in real-world scenarios. Therefore, in this paper, we explore a DRL-based approach, leveraging the fitting capability of neural networks for approximating probability distributions and generating defense policies. We formulate the defense problem for the UAV swarm as a Partially Observable Markov Decision Process (POMDP) for each agent, reflecting the fact that each UAV can only access local and partial observations of the global environment. This formulation can be completely described through its state space, action space, transition probabilities, and reward function as follows.
@@ III-E1 State Space
where $\mathbf{P}(t),\mathbf{V}(t),\mathbf{H}(t)$, and $\mathbf{F}(t)$ denote the positions, velocities, headings, and frequencies of all UAVs at time $t$, respectively. For example, $\mathbf{P}(t)=[\mathbf{p}_{1}(t),\mathbf{p}_{2}(t),\ldots,\mathbf{p}_{N}(t)]$ and $\mathbf{F}(t)=[f_{1}(t),f_{2}(t),\ldots,f_{N}(t)]$. Moreover, $\mathbf{L}(t)$ is the binary indicator of the current leader UAV, which is the UAV that receives commands from the GCS and coordinates the swarm, such that $\mathbf{l}_{i}(t)=1$ when $i=\mathbb{L}^{\mathcal{N}}(t)$, otherwise 00. Finally, $\mathcal{E}(t)$ denotes the communication graph at time $t$, which captures the connectivity between UAVs based on their positions and frequencies (can be derived from Eq. (1)).
@@ III-E2 Action Space
We assume that the agent makes defense decisions in an independent manner, implying that it may decide to take different actions on specific UAVs. At each time step, an agent $i$ can execute one of three MTD mechanisms as described in Section III-C. Hence, we can define the action space for any state in $S_{t}$ as
@@
Specifically, $\tilde{A}_{t}^{\mathbb{L}}=\{\mathbf{l}_{1}(t),\mathbf{l}_{2}(t),...,\mathbf{l}_{N}(t)\}$ denotes the decisions of leader switching at time step $t$. Similarly, $\tilde{A}_{t}^{\mathbb{R}}=\{a_{t,1}^{\mathbb{R}},a_{t,2}^{\mathbb{R}},...,a_{t,N}^{\mathbb{R}}\}$ denotes the actions of route mutation for all UAVs, where $a_{t,n}^{\mathbb{R}}=1$ represents that the $i$-th UAV plays the role of relay node whereas $a_{t,n}^{\mathbb{R}}=-1$ means canceling its role. Furthermore, $\tilde{A}_{t}^{\mathbb{F}}=\{a_{t,1}^{\mathbb{F}},a_{t,2}^{\mathbb{F}},...,a_{t,N}^{\mathbb{F}}\}$ shows a global action of frequency hopping on this UAV swarm, where $\tilde{A}_{t}^{\mathbb{F}}=[1,1,\ldots,1]$ triggers the change of the communication frequency for all UAVs and the GCS. It is worth noting that a zero value indicates no action, that is, this MTD mechanism will not be executed at this time. Therefore, the action selected by agent $i$ at time step $t$ can be represented as
@@ III-E3 Transition Probability
Let $S_{t}\times A_{t}\times S_{t+1}\rightarrow[0,1]$ represent the state transition function. The environment transitions to a new state $s^{\prime}$ according to the transition probability $P(s^{\prime}|s,a)$, where $s$ is the current state at time $t$ and $a\in A_{t}$ is the action selected by all agents. This probability is influenced by the system state, the attacker’s strategy, and the defender’s chosen actions. In the case of deterministic transitions, the probability becomes binary and can be formulated as follows:
@@ III-E4 Reward Function
where $\alpha$, $\beta$, $\zeta$, $\eta$, and $\xi$ are coefficients for balancing the reward. First, the reward of communication connectivity is defined by $R_{\text{com}}^{i}(t)=e_{i}(t)$, which is a binary indicator that represents whether agent $i$ is connected to the leader or the GCS at time step $t$. Second, the reward of formation is defined by $R_{\text{form}}^{i}(t)=1-|\mathbf{p}_{i}(t)-\mathbf{p}_{i}^{*}(t)|/\delta$, which captures the deviation of agent $i$ from its ideal position $\mathbf{p}_{i}^{*}(t)$, normalized by a threshold $\delta$.
Then, we can define the cost of defense actions as $C^{i}(t)=\sum_{i=1}^{N}\mathbf{l}_{i}(t)+a_{t,i}^{\mathbb{R}}+a_{t,i}^{\mathbb{F}}$, which represents the total number of relay actions taken by agent $i$ at time step $t$. The penalty for being under effective attack is defined as $P_{\text{atk}}^{i}(t)=E^{\mathcal{N}}_{i}(t)+\frac{1}{2}E^{\mathcal{L}}_{i\mathbb{L}^{\mathcal{N}}(t)}(t)$, which captures the impact of node and link attacks on agent $i$. Finally, the penalty for excessive velocity is defined as $P_{\text{vel}}^{i}(t)=\frac{v_{i}(t)-v_{\text{pat}}}{v_{\max}-v_{\text{pat}}}$ which penalizes agent $i$ if its speed exceeds the allowed patrol speed.
@@ IV-B Local Policy Optimization
In the local scenario, deep neural network (DNN) models are utilized to build the learning agent. Each agent maintains a policy network $\pi_{\theta_{i}}(a|o)$ parameterized by $\theta_{i}$. At each time step $t$, the agent receives a local observation $o_{t}^{i}$ and samples an action $a_{t}^{i}$ from the policy as
@@
The policy updating procedure adjusts the parameter $\theta_{i}$ to improve the expected long-term cumulative reward of the agent $i$ for action $a_{t}^{i}$ given state $s_{t}^{i}$. The values of actions for sequential observations are measured with the action-value function (or Q-function) to evaluate the expected total return per action. Here, the Q-function is formulated as
@@
where $\Omega_{t}^{i}$ is the expected return for agent $i$ at time step $t$, which can be computed as the discounted sum of future rewards as:
@@
Each agent stores its local experience tuples $(o_{t}^{i},a_{t}^{i},r_{t}^{i},o_{t+1}^{i})$ in a buffer $\mathcal{D}_{i}$. After collecting sufficient experience, the agent performs a policy gradient update. The gradient of the function should be calculated with respect to as follows:
@@
For optimizing the objective function, the policy parameters are updated by minimizing the following loss function:
@@
where the negative sign indicates that we want to maximize the expected return. In addition, we incorporate an entropy regularization term to encourage exploration and prevent premature convergence to suboptimal policies. The modified loss function becomes
@@
where $\hat{R}_{t}^{i}$ is the normalized return, $H(\cdot)$ is the entropy regularization term to encourage exploration, and $\mu$ is the entropy coefficient. In practice, the loss function for each agent can be expressed as

### v2

@@ III-C Moving Target Defense Mechanisms
Figure 2: Overview of the proposed lightweight and coordinated Moving Target Defense (MTD) mechanisms for UAV swarm networks. (a) Leader Switching: Dynamically reassigns the logical leader role among UAVs, thereby enhancing resilience by preventing persistent targeting of a single node/link; (b) Route Mutation: Alters communication paths by selecting alternative relay UAVs, enabling critical messages to bypass compromised links and maintain connectivity under attack; (c) Frequency Hopping: Simultaneously changes the communication frequency channels of all UAVs and the ground control station, increasing uncertainty for attackers and disrupting their ability to sustain effective DoS attacks.
@@
To address these challenges, we propose a suite of lightweight and coordinated MTD mechanisms specifically tailored for UAV swarm networks. These mechanisms are selected for their ability to be executed rapidly, with minimal resource consumption, and without disrupting the primary mission of the UAV swarm. The goal is to create a moving target for attackers, making it difficult for them to maintain effective attacks or predict the system’s defensive responses. Fig. 2 provides an overview of the specific MTD-based mitigation mechanisms adopted in this work.
@@ III-C1 Leader Switching
Leader switching designates a new UAV as the swarm leader when the current leader is persistently targeted or isolated by attacks. It is important to note that leader switching is a logical (virtual) role reassignment that can be executed rapidly and with minimal overhead and does not require any physical movement within the swarm. This mechanism only involves updating the logical leader role in the existing control framework, requiring negligible computational resources and no additional communication overhead, thus having minimal impact on the physical formation or flight of the UAVs.
Let $\mathbb{L}^{\mathcal{N}}(t)$ denote the swarm leader at time $t$. The system can switch to a new leader chosen from the set of eligible UAVs after the leader switching occurs, such that
@@
Here, $i$ cannot be 0 (the GCS) or the current leader and $\tau^{\mathbb{L}}_{\text{exec}}$ denotes the execution time that the leader switching mechanism takes. After the switch, the new leader takes over the responsibility of receiving control commands from the GCS and disseminating them to the follower UAVs through the flying Ad-Hoc network (FANET).
@@ III-C2 Route Mutation
Usually, the command is transferred from the GCS to the leader, and is then forwarded to the follower UAVs. However, when a direct communication link is blocked by an attack, we exploit route mutation that selects an alternative relay UAV to forward messages. For example, when the communication link between the current leader $\mathbb{L}^{\mathcal{N}}(t)=i$ and a follower UAV $j$ is under attack, then the relay UAV $r\in\mathcal{V}\setminus\{0,i,j\}$ should satisfy the following rules
@@
This operation dynamically reconfigures the communication graph in real time, allowing the command to be forwarded through the relay UAV and successfully delivered to the follower, thus mitigating the impact of the targeted link attack, i.e., $E^{\mathcal{L}}_{ij}(t+\tau^{\mathbb{R}}_{\text{exec}})=0$, where $\tau^{\mathbb{R}}_{\text{exec}}$ is the time for updating the routing table.
By leveraging existing routing protocols and requiring only simple updates, this mechanism is lightweight, relying solely on local computations without introducing significant communication overhead. When there is no ongoing attack or when the attack fails to produce a substantial impact, the system can further optimize network performance by restoring the relay-based communication path back to a direct link. This reduces communication latency and improves overall efficiency without compromising the defense capability.
@@ III-C3 Frequency Hopping
In practical UAV swarm deployments, frequency hopping can be implemented in a lightweight and coordinated manner. Each UAV and the ground control station maintain a synchronized hopping sequence, which can be generated using a pseudo-random algorithm or a shared secret key and be performed rapidly to minimize communication disruption against persistent and sophisticated DoS attacks.
@@ III-D Problem Formulation
For the UAV swarm network, its objective is to defend against the malicious DoS attacker. The UAVs need to choose the proper defensive actions to optimize their performance and cost in a distributed setting. Therefore, we formulate the optimization problem as
@@
where $|\mathcal{E}(t)|$ is the cardinality of this set and $C^{i}(t)$ is the defense cost of UAV $i$ at each time slot. Intuitively, the multi-objective optimization problem (P1) has the following objectives: (i) The first part aims to minimize the total number of effective node and link attacks over time, which is represented by the sum of $E^{\mathcal{N}}_{i}(t)$ and $E^{\mathcal{L}}_{ij}(t)$. (ii) The second part aims to maximize the network connectivity, which is represented by the maximum number of active communication links $|\mathcal{E}(t)|$ at any time step. (iii) The third part aims to minimize the deviation of each UAV from its ideal formation position, which is represented by the sum of the distance between each UAV’s position and its ideal position $\mathbf{p}_{i}^{*}(t)$. (iv) Finally, the fourth part aims to minimize the cumulative defense cost, which is represented by the sum of defense costs $C^{i}(t)$ for all UAVs.
The constraints ensure that the UAV swarm operates within the defined parameters and have the following practical meanings: (i) Constraints C1 enforce physical and communication limits, such as minimum separation between UAVs, attack effectiveness, and valid communication links. (ii) Constraint C2 ensures that the heading angle of each UAV is within the range $0,2\pi)$, reflecting feasible orientation in 2D space. (iii) Constraint C3 restricts the speed of each UAV to be within the patrol speed $r\omega$ and the maximum speed $v_{\max}$, ensuring safe and energy-efficient flight. (iv) Constraint C4 ensures that all UAVs operate on the same frequency channel at any time step, which is essential for maintaining communication connectivity. (v) Finally, constraint C5 ensures realistic attack cycles by stipulating that the total duration of any number of attacks remains within the mission time, where the effective duration of a single attack round $\tau_{\text{eff}}$ is the sum of the attack period and the reconnaissance period.
@@ III-E Optimization Analysis
It is worth noting that the multi-objective optimization problem above might not be solved by using traditional methods since obtaining accurate knowledge of the system dynamics and the attacker’s behavior is impractical in real-world scenarios. Therefore, in this paper, we explore a DRL-based approach, leveraging the fitting capability of neural networks for approximating probability distributions and generating defense policies.
To formalize the defense decision problem, we adopt a Partially Observable Markov Decision Process (POMDP) framework. A POMDP is particularly suitable for our UAV defense scenario because: (i) each UAV can only access local information rather than the complete system state, (ii) the environment changes in response to both defense and attack actions, and (iii) we need sequential decision-making under uncertainty. This formulation can be completely described through its state space, action space, transition probabilities, and reward function as follows.
@@ III-E1 State Space
where $\mathbf{P}(t),\mathbf{V}(t),\mathbf{H}(t)$, and $\mathbf{F}(t)$ denote the positions, velocities, headings, and frequencies of all UAVs at time $t$, respectively. For example, $\mathbf{P}(t)=[\mathbf{p}_{1}(t),\mathbf{p}_{2}(t),\ldots,\mathbf{p}_{N}(t)]$ and $\mathbf{F}(t)=[f_{1}(t),f_{2}(t),\ldots,f_{N}(t)]$. Moreover, $\mathbf{L}(t)$ is the binary indicator of the current leader UAV, which is the UAV that receives commands from the GCS and coordinates the swarm, such that $\mathbf{l}_{i}(t)=1$ when $i=\mathbb{L}^{\mathcal{N}}(t)$, otherwise 0. Finally, $\mathcal{E}(t)$ denotes the communication graph at time $t$, which captures the connectivity between UAVs based on their positions and frequencies (can be derived from Eq. (1)).
@@ III-E2 Action Space
At each decision point, each UAV agent needs to choose a defensive action. The set of all possible actions for an agent forms its action space. We assume that the agent makes defense decisions in an independent manner, implying that it may decide to take different actions on specific UAVs. At each time step, an agent $i$ can execute one of three MTD mechanisms as described in Section III-C. Hence, we can define the action space for any state in $S_{t}$ as
@@
Specifically, $\tilde{A}_{t}^{\mathbb{L}}=\{\mathbf{l}_{1}(t),\mathbf{l}_{2}(t),...,\mathbf{l}_{N}(t)\}$ denotes the decisions of leader switching at time step $t$. Similarly, $\tilde{A}_{t}^{\mathbb{R}}=\{a_{t,1}^{\mathbb{R}},a_{t,2}^{\mathbb{R}},...,a_{t,N}^{\mathbb{R}}\}$ denotes the actions of route mutation for all UAVs, where $a_{t,n}^{\mathbb{R}}=1$ represents that the $i$-th UAV plays the role of relay node whereas $a_{t,n}^{\mathbb{R}}=-1$ means canceling its role. Although the action space is formally defined for all agents, the initiation of global MTD mechanism (i.e., frequency hopping) is a privileged operation executed only by the current leader agent $\mathbb{L}^{\mathcal{N}}(t)$. When the leader’s policy selects frequency hopping, the corresponding command is broadcast to all follower agents for synchronized execution. We define it as $\tilde{A}_{t}^{\mathbb{F}}=\{a_{t,1}^{\mathbb{F}},a_{t,2}^{\mathbb{F}},...,a_{t,N}^{\mathbb{F}}\}$, where $\tilde{A}_{t}^{\mathbb{F}}=[1,1,\ldots,1]$ triggers the change of the communication frequency for all UAVs and the GCS. This ensures that while all agents learn a comprehensive policy, only the leader’s decision on global matters is enacted, thus resolving the coordination challenge.
Furthermore, It is worth noting that a zero value indicates no action, that is, this MTD mechanism will not be executed at this time. Therefore, the action selected by agent $i$ at time step $t$ can be represented as
@@ III-E3 Transition Probability
Let $S_{t}\times A_{t}\times S_{t+1}\rightarrow[0,1]$ represent the state transition function. The environment transitions to a new state $s^{\prime}$ according to the transition probability $P(s^{\prime}|s,a)$, where $s$ is the current state at time $t$ and $a\in A_{t}$ is the action selected by all agents. This probability is influenced by the system state, the attacker’s strategy, and the defender’s chosen actions. In the case of deterministic transitions, the probability becomes binary and can be formulated as
@@ III-E4 Reward Function
where $\alpha,\beta,\zeta,\eta$, and $\xi$ are weighting coefficients. These coefficients are crucial for balancing the trade-offs between mission objectives (e.g., connectivity, formation) and operational costs (e.g., defense actions, attack penalties). Their final values were determined through an empirical tuning process, as detailed in Section V-A. The components are defined as follows:
First, the reward of communication connectivity is defined by $R_{\text{com}}^{i}(t)=e_{i}(t)$, which is a binary indicator (defined in Eq. (13)) that represents whether agent $i$ is connected to the leader or the GCS at time step $t$. Second, the reward of formation is defined by $R_{\text{form}}^{i}(t)=1-|\mathbf{p}_{i}(t)-\mathbf{p}_{i}^{*}(t)|/\delta$, which captures the deviation of agent $i$ from its ideal position $\mathbf{p}_{i}^{*}(t)$, normalized by the maximum allowed deviation threshold $\delta$ (defined in Section III-A).
Then, we define the cost of defense actions for agent $i$ as $C^{i}(t)=\sum_{j=1}^{N}\mathbf{l}_{j}(t)+a_{t,i}^{\mathbb{R}}+a_{t,i}^{\mathbb{F}}$. This cost function is composed of three parts: (i) The first term $\sum_{j=1}^{N}\mathbf{l}_{j}(t)$ represents the cost associated with the leader switching action. Since leader switching is a global action that affects the entire swarm’s coordination, its cost is accounted for across all agents. (ii) The terms $a_{t,i}^{\mathbb{R}}$ and $a_{t,i}^{\mathbb{F}}$ are binary indicators (defined in Eq. (15)) representing whether agent $i$ executes a route mutation or participates in a frequency hop, respectively. The penalty for being under effective attack is defined as $P_{\text{atk}}^{i}(t)=E^{\mathcal{N}}_{i}(t)+\frac{1}{2}E^{\mathcal{L}}_{i\mathbb{L}^{\mathcal{N}}(t)}(t)$, which captures the impact of node and link attacks on agent $i$. Finally, the penalty for excessive velocity is defined as $P_{\text{vel}}^{i}(t)=\frac{v_{i}(t)-v_{\text{pat}}}{v_{\max}-v_{\text{pat}}}$ which penalizes agent $i$ if its speed exceeds the allowed patrol speed.
The reward function effectively transforms the multi-objective optimization problem P1 into a single-objective learning goal. By structuring the reward as a weighted sum of terms corresponding to the objectives in P1 where maximization goals are positive rewards and minimization goals are penalties, we convert the problem into maximizing a single cumulative reward. The coefficients control the trade-offs between competing goals like maximizing security and minimizing cost.
@@ IV-B Local Policy Optimization
We select a policy gradient (PG) based approach for local policy optimization due to its specific advantages in addressing the UAV defense problem. First, PG methods naturally learn a stochastic policy, which aligns perfectly with the MTD principle by introducing unpredictability into the UAV swarm. This makes it significantly harder for an attacker to counter the defensive actions. Second, PG methods are well-suited for the complex, multi-dimensional action space of our problem and generally offer more stable convergence properties compared to value-based methods like DQN, which is crucial for learning a reliable policy in our dynamic environment.
In detail, deep neural network (DNN) models are utilized to build the learning agent in the loacl scenario. Each agent maintains a policy network $\pi_{\theta_{i}}(a|o)$ parameterized by $\theta_{i}$. At each time step $t$, the agent receives a local observation $o_{t}^{i}$ and samples an action $a_{t}^{i}$ from the policy as
@@
The policy updating procedure adjusts the parameter $\theta_{i}$ to improve the expected long-term cumulative reward of the agent $i$ for action $a_{t}^{i}$ given state $s_{t}^{i}$. To evaluate how good an action is in a given state, we use an action-value function (or Q-function), which estimates the expected total return per action. The Q-function is formulated as
@@
where $\Omega_{t}^{i}$ is the expected return for agent $i$ at time step $t$, which can be computed as the discounted sum of future rewards as
@@
Each agent stores its local experience tuples $(o_{t}^{i},a_{t}^{i},r_{t}^{i},o_{t+1}^{i})$ in a buffer $\mathcal{D}_{i}$. After collecting sufficient experience, the agent performs a policy gradient update, which adjusts the policy network’s parameters $\theta_{i}$ in the direction that increases the expected reward. The gradient of the function should be calculated with respect to as follows
@@
For optimizing the objective function, the policy parameters are updated by minimizing the following loss function as
@@
where the negative sign indicates that we want to maximize the expected return. To encourage exploration and prevent the agent from getting stuck in a suboptimal strategy, we add an entropy regularization term. The modified loss function becomes
@@
where $\hat{R}_{t}^{i}$ is the normalized return, and $H(\cdot)$ is the entropy regularization term to encourage exploration. The entropy coefficient $\mu$ controls the trade-off between exploiting the current best-known actions and exploring new ones to discover potentially better strategies. A higher $\mu$ encourages more randomness in action selection, while a lower value leads to a more deterministic policy. Its value was determined empirically, as detailed in Section V-A.
In practice, the loss function can be expressed as